Find the quadratic equation If r1+r2 = 0 , r1*r2 = 6
Answer by richwmiller(17219)

for clarity let r1=r and s=r2
r+s=0
r*s=6
(r-i*sqrt(6))*(r+i*sqrt(6))*(s-i*sqrt(6))*(s+i*sqrt(6))=0
(r^2+6)*(s^2+6)=0
r^2s^2+6r^2+6s^2+36=0